OpenStudy (anonymous):

Choose the simplified form of the radical expression. 4 sqrt 8 / sqrt 2 8 16 2 sqrt 16 4 sqrt 16

OpenStudy (anonymous):

\[\frac{ 4\sqrt{8} }{ \sqrt{2} }=\frac{ 4\sqrt{4\times2} }{ \sqrt{2} }=\frac{ 4\times2\sqrt{2} }{2 }=4\times2=8\]